OpenStudy (anonymous):

1 + cos(6x) = _____?

OpenStudy (sirm3d):

hint: \[\cos 2A=2\cos^2A-1\] or after transposition, \[1+\cos 2A=2\cos^2A\]

OpenStudy (anonymous):

2cos2(3x) ?

OpenStudy (anonymous):

@sirm3d

OpenStudy (sirm3d):

yup